OpenStudy (anonymous):

Given the function f(x) = 2(x + 8), find x if f(x) = 12 2 40 −2 14

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

12 = 2(x+8) 12 = 2x + 16 -4 = 2x -2 = x

jhonyy9 (jhonyy9):

2x+16=12 2x=-4 x=-2